Write this in standard form:
the line perpendicular to -2x + 3y = 9 through (-1,-3)

Answer by stanbon(75887)

Write this in standard form:
the line perpendicular to -2x + 3y = 9 through (-1,-3)
------
Given line::
y = (2/3)x + 3
slope = 2/3
------
Perpendicular line::
slope = -3/2
Form:: y = mx + b
Solve for "b":
-3 = (-3/2)(-1) = b
(-6/2)-(3/2) = b
-9/2
------
Equation:
y = (-3/2)x - (9/2)
-------------------------
Cheers,
Stan H.